### v4

# Easy3D-Labels: Supervising Semantic Occupancy Estimation with 3D Pseudo-Labels for Automotive Perception

Abstract
In perception for automated vehicles, safety is critical not only for the driver but also for other agents in the scene, particularly vulnerable road users such as pedestrians and cyclists. Previous representation methods, such as Bird’s Eye View, collapse vertical information, leading to ambiguity in 3D object localisation and limiting accurate understanding of the environment for downstream tasks such as motion planning and scene forecasting. In contrast, semantic occupancy provides a full 3D representation of the surroundings, addressing these limitations. Furthermore, self-supervised semantic occupancy has seen increased attention in the automated vehicle domain. Unlike supervised methods that rely on manually annotated data, these approaches use 2D pseudo-labels, improving scalability by reducing the need for labour-intensive annotation. Consequently, such models employ techniques such as novel view synthesis, cross-view rendering, and depth estimation to allow for model supervision against the 2D labels. However, such approaches often incur high computational and memory costs during training, especially for novel view synthesis. To address these issues, we propose Easy3D-Labels, which are 3D pseudo-ground-truth labels generated using Grounded-SAM and Metric3Dv2, with temporal aggregation for densification, permitting supervision directly in 3D space. Easy3D-Labels can be readily integrated into existing models to provide model supervision, yielding substantial performance gains, with mIoU increasing by 45% and RayIoU by 49% when applied to OccNeRF on the Occ3D-nuScenes dataset. Additionally, we introduce EasyOcc, a streamlined model trained solely on these 3D pseudo-labels, avoiding the need for complex rendering strategies, and achieving 15.7 mIoU on Occ3D-nuScenes. Easy3D-Labels improve scene understanding by reducing object duplication and enhancing depth estimation accuracy, as reflected by improvements in the RayIoU metric. These findings highlight the importance of foundation models, temporal information, and 3D loss formulation in self-supervised learning for comprehensive scene understanding. Our Easy3D-Labels are available open-source on Mendeley
Figure 1: Easy3D-Labels generation: We project semantic labels into 3D using depth maps and employ temporal aggregation and object filtering for enhanced label quality.
I INTRODUCTION
Safety in the domain of automated vehicle perception is critical to establish trust for both drivers and other road users [40]. In 2021, the global annual road traffic deaths were estimated at 1.19 million, with pedestrians accounting for 21% of these fatalities [50]. It is evident that vulnerable road users account for a large proportion of road injuries and deaths. This emphasises the need for accurate and timely detection of agents in the scene to enable the vehicle to predict trajectories and make decisions in high-risk situations.
Detection accuracy is closely tied to how the scene is represented, as restrictive representations can limit both perceptual coverage [30] and the completeness of spatial reasoning [32]. Following recent progress in machine learning, many advanced approaches for automated vehicle perception have been developed [52, 12]. One prominent approach is semantic occupancy estimation [49], whose discretised 3D representation enables more flexible scene modelling compared to earlier methods such as Bird’s Eye View and 3D bounding boxes that contain the aforementioned limitations.
Specifically, self-supervised semantic occupancy estimation is especially advantageous, largely due to its reduced reliance on manually annotated occupancy labels, offering improved scalability compared to supervised approaches [17]. However, these models still depend on labels in the 2D image space, using Vision Language Models (VLM) [27, 36] and Visual Foundation Models (VFM) [22, 14] to address semantic and depth ambiguities in the absence of ground-truth 3D labels. To enable supervision in 2D, models typically use novel view synthesis techniques to render the 3D scene in image space, employing NeRF-based volume rendering [33, 55] or 3D Gaussian Splatting [21, 9, 19]. However, relying on 2D rendering introduces high computational cost and leads to limitations such as bias towards nearby objects [12], inaccurate detections, and object duplication [41].
To address these issues, we propose a 3D labelling technique, Easy3D-Labels, which generates labels by projecting Grounded-SAM [37] 2D semantics into 3D space using Metric3Dv2 [14] depth maps, illustrated in Figure 1. This allows the model to jointly learn semantics and spatial geometry through a single pseudo-loss function. Furthermore, we aggregate temporal samples, limited to static objects, to avoid duplicating dynamic objects to increase density. Prior work has shown the value of temporal information for improving performance [44, 57, 25]. The aggregation of temporal data in the supervision pipeline reduces additional computational overhead, as it is typically performed at inference time. Easy3D-Labels provide several important advantages. Firstly, the removal of the need for novel view synthesis and depth estimation during training reduces computational cost. Secondly, they enable efficient aggregation of temporal information, which is important for spatial understanding. Lastly, they support more comprehensive scene understanding in both visible and occluded regions
Furthermore, Easy3D-Labels can be easily added to existing models as an auxiliary loss to boost performance. We explore the integration of these labels in three previous models: SelfOcc, OccNeRF, and GaussianOcc. As recent advances often develop in isolation, we present a complementary method to enhance model compatibility and generalization. Additionally, we introduce EasyOcc, a streamlined model that solely uses Easy3D-Labels for loss computation, demonstrating that complex rendering techniques are not necessary for noteworthy model performance. Furthermore, our model requires no LiDAR supervision or deployment of foundation models at inference.
In summary, our main contributions are as follows:
  • Easy3D-Labels: We introduce an approach that leverages Grounded-SAM and Metric3Dv2 to generate 3D pseudo-labels for loss computation directly in 3D space. Our 3D pseudo-labels can be effortlessly integrated into existing models via an auxiliary loss function, yielding improvements of 45% in mIoU.
  • Segmentation of Dynamic Classes: Accurate segmentation of dynamic classes is critical for safety in autonomous perception. In the case of SelfOcc, incorporating our 3D pseudo-labels improves pedestrian segmentation performance by over 600%.
  • Holistic Scene Representation: Our labels enable a more comprehensive scene representation, resulting in a 49% increase in RayIoU for OccNeRF.
This paper is structured as follows: Section II reviews related work, Sections III and IV describe Easy3D-Labels and the EasyOcc model, Section V presents results and ablations, and finally Section VI concludes the paper.
II LITERATURE REVIEW
In Subsection II-A, we review semantic occupancy models, and in Subsection II-B, we discuss prior use of 3D pseudo-labels.
II-A SEMANTIC OCCUPANCY ESTIMATION
In perception for automated vehicles, Bird’s Eye View (BEV) methods have historically been dominant due to their simple yet effective scene representation [11, 34, 24, 28]. Recently, semantic occupancy estimation has gained attention, driven by benchmark datasets [49, 45, 64, 46] with accurate annotations, generated from manually labeled LiDAR data from the nuScenes automated vehicles dataset [6]. This shift led to the creation of supervised occupancy estimation models [43, 38], with improvements from techniques such as Gaussian Splatting [18, 16, 59, 39], multi-modal fusion [13, 53, 31], and object deduplication [41]. Following this, self-supervised counterparts of these models emerged, particularly due to their flexibility in training strategy, requiring no manually annotated ground truth labels. In this study, we modify three self-supervised models, SelfOcc, OccNeRF, and GaussianOcc, which all follow a common pipeline of 2D image encoding, 3D feature lifting, and 3D voxel refinement prior to rendering.
SelfOcc employs an MLP to predict signed distance field (SDF) values, color, and semantic features from the 3D volume, for rendering depth, color, and semantics [17]. Depth supports multi-frame photometric consistency, color for comparison to the RGB image, and semantics against 2D pseudo-labels from OpenSeeD [58]. Semantics and occupancy are both computed via the SDF, with both contributing to the final scene representation. In our implementation, pseudo-loss is applied exclusively to the semantic voxel.
OccNeRF deploys NeRF-style volume rendering to render both depth and semantic information, with depth supporting multi-frame photometric consistency, while semantics are compared against pseudo-labels from Grounded-SAM [37]. GaussianOcc builds on OccNeRF with Gaussian rasterization [21] for rendering both semantics and depth [9]. For use in multi-frame consistency, it estimates pose transformations using a 6D pose network instead of ground-truth poses, which is more effective, given the nuScenes dataset’s lack of $z$-axis translation in ego-vehicle transformations [6]. For both methods, pseudo-loss is applied to the semantic voxel grid.
Other state-of-the-art methods employ Gaussian scene representations, which prove beneficial for reduced memory consumption due to their sparse nature. Methods employ techniques including self-attention and image cross-attention mechanisms [19, 5], foundation models during inference [19, 57], and temporal flow modelling [5], with supervision labels originating again from foundation models [47, 14, 37, 2, 20]. AutoOcc is an estimation and labelling pipeline which uses numerous attention mechanisms [8, 26, 1] and foundation models [35, 54, 60] in a test-time manner. However, this method does not explore the deployment of its labels into existing models for enhancement.
Our proposed model, EasyOcc, shares a similar pipeline with GaussianOcc. However, it omits several components: pose estimation, novel view synthesis, and multi-frame photometric consistency. Instead, EasyOcc solely leverages Easy3D-Labels for supervision. Despite its simplified design, EasyOcc outperforms other models that employ complex training paradigms in the mIoU and RayIoU, as detailed in Section V.
II-B PSEUDO-LABELS
The use of 2D pseudo-labels in self-supervised semantic occupancy models has been extensively studied, particularly through the application of VLMs [37, 36, 2] and VFMs [35, 14]. These models have demonstrated utility across various domains, including medical applications [56, 29, 51, 10], and robotics [15, 7]. In perception for automated vehicles, they can play a crucial role in addressing the challenge of missing ground-truth labels, particularly in resolving semantic and depth ambiguities. VLMs help mitigate semantic ambiguity by leveraging both spatial and linguistic cues to produce pixel-level semantic maps. Depth ambiguity, while partially addressed using multi-frame photometric consistency, is significantly reduced through the use of metric depth VFMs [14, 35, 4], which provide accurate pixel-level depth maps, shown to increase model performance substantially [19, 5]. Semantic maps and depth maps serve as supervision signals against renders of the semantic voxel grid. Nonetheless, a noticeable performance gap persists between supervised and self-supervised models, indicating a need for a more nuanced integration of foundation models, potentially by focusing on the dimensionality of the labels.
More recently, the use of 3D pseudo-labels has gained attention, which aligns with the space in which final estimations are made. A notable example is AGO [23], which combines Grounded-SAM [37] with LiDAR point cloud data, utilizing multi-frame aggregation, point cloud ray casting, and semantic voting to generate richer labels for training. However, this approach necessitates equipping the vehicle with a LiDAR sensor, which introduces high cost and complexity, including the need for careful synchronization with the camera system. Additionally, the output generated by the aforementioned AutoOcc can be considered a form of 3D pseudo-labels.
Our proposed method, Easy3D-Labels, does not rely on LiDAR and instead combines Metric3Dv2 [14] for depth estimation with Grounded-SAM [37] for semantic segmentation to generate 3D pseudo-ground-truth labels. These labels are further refined through outlier removal, occupancy thresholding, and temporal aggregation, enabling models to learn a more holistic scene representation.
III METHODOLOGY I: EASY3D-LABELS
This section outlines the generation of Easy3D-Labels in Subsection III-A, followed by an evaluation of their quality in Subsection III-B, including comparisons with ground truth.
III-A GENERATION OF EASY3D-LABELS
This section presents the key contribution of this paper: the generation of our 3D pseudo-labels from semantic maps of Grounded-SAM [37] and depth maps from Metric3Dv2 [14]. This method enables loss computation directly in 3D voxel space, eliminating the need for view synthesis and aligning our approach with supervised training pipelines. The process is illustrated in Figure 1 and divided into three steps: semantic point cloud generation (Subsubsection III-A1), densification (Subsubsection III-A2), and voxelization (Subsubsection III-A3).
III-A1 SEMANTIC POINT CLOUD GENERATION
This stage will detail the generation of a semantic point cloud for an arbitrary sample. Semantic maps are sourced from the OccNeRF repository [55], which are generated using Grounded-SAM, while depth maps are generated with the Giant variant of Metric3Dv2 for optimal training performance [14]. For each camera, $i\in\{1,\dots,6\}$, in a sample, given the corresponding semantic map $S_{i}\in\mathbb{R}^{H\times W}$, depth map $D_{i}\in\mathbb{R}^{H\times W}$, camera intrinsic matrix $K_{i}\in\mathbb{R}^{3\times 3}$, and camera-to-global transformation $\mathbf{T}_{\text{camera, }i}^{\text{global}}\in\mathbb{R}^{4\times 4}$, each arbitraty pixel $(u,v)\in[0,W)\times[0,H)$ is projected into the dehomogenised 3D global coordinates in Equation (1):
$\mathbf{P}_{\text{global, }i}^{(u,v)}=\mathbf{T}_{\text{camera, }i}^{\text{global}}\begin{bmatrix}D_{i}(u,v)\cdot K_{i}^{-1}\begin{bmatrix}u\\ v\\ 1\end{bmatrix}\\ 1\end{bmatrix}$ (1)
Following this, each projected pixel, $\mathbf{P}_{\text{global,}i}^{(u,v)}$, is then decorated with its corresponding semantic pixel $S_{i}(u,v)$, to yield a semantic point cloud, $\mathcal{P}^{\text{sem}}_{i}$, seen in Equation (2), where $\mathcal{L}=\{0,1,\dots,17\}$ denotes the semantic label space.
$\mathcal{P}^{\text{sem}}_{i}=\left\{\left(\mathbf{P}_{\text{global, }i}^{(u,v)},\;S_{i}(u,v)\right)\right\},\quad\mathcal{P}^{\text{sem}}_{i}\subset\mathbb{R}^{3}\times\mathcal{L}$ (2)
Finally, we aggregate the semantic point cloud for each camera, $\mathcal{P}^{\text{sem}}_{i}$, into a unified semantic point cloud, $\mathcal{P}^{\text{sem}}$, expressed in Equation (3).
$\mathcal{P}^{\text{sem}}=\bigcup_{i=1}^{N}\left\{\mathcal{P}^{\text{sem}}_{i}\right\}$ (3)
To improve spatial accuracy, outlier removal is performed on $\mathcal{P}^{\text{sem}}$ using the Open3D library [62], yielding a consolidated semantic point cloud in the global coordinate frame.
III-A2 SEMANTIC POINT CLOUD DENSIFICATION
Following the previous step, point clouds for each sample, $\mathcal{P}_{T}^{\text{sem}}$, are densified using temporal semantic point clouds, $\mathcal{P}_{T{-}\Delta}^{\text{sem}}$, for all $\Delta\in\{1,\dots,13\}$. First, we remove dynamic points (e.g., vehicles and pedestrians) in $\mathcal{P}_{T{-}\Delta}^{\text{sem}}$, to prevent object duplication, while retaining static points, such as the sidewalk and drivable surface. Following this, we transform the unions of $\mathcal{P}_{T}^{\text{sem}}$ and $\mathcal{P}_{T{-}\Delta}^{\text{sem}}$ from global coordinates to ego-vehicle coordinates of time $T$ with $\mathbf{T}_{\text{global}}^{\text{ego}}\in\mathbb{R}^{4\times 4}$, which yields the densified semantic point cloud, $\mathcal{P}_{T}^{\text{final sem}}$, in Equation (4):
$\mathcal{P}_{T}^{\text{final sem}}=\mathbf{T}_{\text{global}}^{\text{ego}}\circ\left(\mathcal{P}_{T}^{\text{sem}}\cup\bigcup_{\Delta=1}^{13}\mathcal{P}_{T{-}\Delta}^{\text{sem}}\right)$ (4)
This process ensures that the final voxelisation of $\mathcal{P}_{T}^{\text{final sem}}$ yields labels more closely resembling ground-truth data. The effectiveness of this step is demonstrated in the following Subsection III-B1 and in the ablation study on the EasyOcc model in Subsection V-E2.
III-A3 SEMANTIC POINT CLOUD VOXELIZATION
Once $\mathcal{P}_{T}^{\text{sem}}$ is obtained, it is voxelized to obtain the final 3D pseudo-label. The bounds are defined by the Occ3D-nuScenes [45] ground truth: $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, using a voxel resolution of $0.4\text{m}^{3}$, expressed in the ego-frame coordinate system of the current sample. Given the high density of $\mathcal{P}_{T}^{\text{sem}}$ due to the aggregation of many temporal samples, a voxel is considered occupied only if it contains a minimum of ten points; otherwise, it is treated as empty. This threshold helps mitigate the influence of stray points that could otherwise result in erroneous voxelization. For voxels classified as occupied, the semantic label is assigned based on the majority class among the contained points.
III-B EASY3D-LABELS QUALITY
In this section, we compare our Easy3D-Labels with the ground-truth labels from Occ3D-nuScenes [6], both quantitatively and qualitatively.
III-B1 QUANTITATIVE ANALYSIS
In Figure 2, we compare Easy3D-Labels using varying numbers of aggregated temporal samples against the Occ3D-nuScenes ground-truth labels. The result follows a logarithmic trend, indicating saturation, where aggregating more temporal samples provides diminishing returns in mIoU. The optimal number of temporal samples is found to be 13, at which point we achieve the highest mIoU score of 15.4. These findings show that incorporating temporal samples improves the similarity of the pseudo-labels to the Occ3D ground truth. The maximum number of temporal samples is capped at 13 due to memory constraints.
Figure 2: Temporal sample aggregation: Easy3D-Labels compared to Occ3D-nuScenes [45] labels for various numbers of aggregated samples.
In Figure 3, we compare our labels generated using different occupancy threshold values, ranging from 1 to 25. The highest performance is observed at a threshold of 3, indicating that even noisy points contribute useful information. For our experiments, we selected a threshold of 10 to balance slightly faster generation time with comparable accuracy.
Figure 3: Occupancy threshold: Easy3D-Labels compared to Occ3D-nuScenes [45] labels for various threshold values in the generation processes.
III-B2 QUALITATIVE ANALYSIS
In Figure 4, we compare four Easy3D-Labels training samples with their corresponding ground-truth labels. The pseudo-labels closely match the ground truth, accurately identifying key scene elements such as roads, vegetation, and buildings. Despite mitigation efforts such as outlier removal and occupancy thresholding, there exist incorrectly labelled voxels primarily due to noise in the depth maps. However, as will be discussed in Subsection V-F, the model’s predicted outputs often appear smoother and more continuous than the ground-truth labels. Densification is lacking in the rightmost sample due to the sample being early in the sequence, resulting in limited aggregation of temporal data.
Figure 4: Pseudo-label comparison: Occ3D-nuScenes [45] ground truth (top) and our Easy3D-Labels (bottom)
IV METHODOLOGY II: MODELS
This section introduces the pseudo-loss, $\mathcal{L}_{\text{Pseudo}}$, in Subsection IV-A, model modifications in Subsection IV-B, and our model, EasyOcc, in Subsection IV-C.
TABLE I: Model configurations: Rendering time denotes the time to render semantics, depth, or features per sample during training, while training time is reported per epoch; (+) indicates additional time from pseudo-loss. * OccNeRF parameters include the NeRF rendering module (training only). ** GaussianOcc parameters include the pose estimation module (training only).
Method Backbone Model Parameters Image Size Epochs Rendering Time Training Time
[…]
IV-A PSEUDO LOSS
Model supervision using Easy3D-Labels will be facilitated by the pseudo-loss function, $\mathcal{L}_{\text{Pseudo}}$. This loss consists of two distinct terms, as shown in (5), where $\lambda$ is a constant initialized at the start of training (an ablation on the value of $\lambda$ is discussed in Table VII). The formulation is adapted from GaussianOcc, where it was initially used as an optional component for training with ground-truth labels [9].
$\mathcal{L}_{\text{Pseudo}}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{Geomety}}$ (5)
As shown in Equation (6), geometry loss is composed of three separate losses: geometric scale loss, semantic scale loss, and Lovász softmax loss [3].
$\mathcal{L}_{\text{Geometry}}=\mathcal{L}_{\text{geom\_scal}}+\mathcal{L}_{\text{sem\_scal}}+\mathcal{L}_{\text{Lovász}}$ (6)
All losses, including cross-entropy, are standard in semantic occupancy estimation, as they effectively penalize misclassifications and support class re-weighting to address dataset imbalance.
IV-B MODIFICATIONS TO EXISTING MODELS
Incorporating $\mathcal{L}_{\text{Pseudo}}$ into the three selected architectures requires considering how the loss function interacts with existing losses and also specific implementation details, as outlined below.
SelfOcc: As described in Section II, SelfOcc predicts semantic occupancy using a two-step process: binary occupancy prediction (occocc), followed by a semantic voxel prediction (semsem), which together produce the final output. This structure introduces two key considerations: (1) a voxel may be classified as occupied by occocc but empty by semsem, resulting in it being considered unoccupied, and (2) a voxel may be classified as unoccupied by occocc but occupied by semsem, leading to it remaining unoccupied.
Through preliminary testing, we observe that excluding occocc from pseudo-loss computation and from the final scene representation improved performance in both IoU and mIoU metrics. This is perhaps explained by the considerations discussed above. Hence, the final scene representation is the semantic voxel, semsem. The final loss function is defined in Equation (7). The additional losses present aid in SDF stability, multi-frame photometric consistency, RGB rendering, and 2D semantic loss.
$\mathcal{L}_{\text{SelfOcc}}=\mathcal{L}_{\text{regularisation}}+\mathcal{L}_{\text{reprojection}}+\mathcal{L}_{\text{rgb}}+\mathcal{L}_{\text{sem}}+\mathcal{L}_{\text{Pseudo}}$ (7)
OccNeRF: Here, we implement pseudo-loss alongside additional loss components to form the final loss function, as shown in Equation (8). The pseudo-loss serves as a complement to the three existing losses in the original OccNeRF model: $\mathcal{L}_{\text{regularisation}}$, $\mathcal{L}_{\text{reprojection}}$, $\mathcal{L}_{\text{sem}}$. These losses regulate voxel occupancy stability, multi-frame photometric consistency, and 2D semantic loss, respectively.
$\mathcal{L}_{\text{OccNeRF}}=\mathcal{L}_{\text{regularisation}}+\mathcal{L}_{\text{reprojection}}+\mathcal{L}_{\text{sem}}+\mathcal{L}_{\text{Pseudo}}$ (8)
GaussianOcc: Given the similarity between GaussianOcc and OccNeRF, our pseudo-loss implementation follows the same approach in GaussianOcc, with one key difference: we omit $\mathcal{L}_{\text{sem}}$ due to sporadic NaN gradients in the convolutional layers of the image encoder during training. The cause of this issue remains unknown. The resulting loss function is defined in Equation (9).
$\mathcal{L}_{\text{GaussianOcc}}=\mathcal{L}_{\text{regularisation}}+\mathcal{L}_{\text{reprojection}}+\mathcal{L}_{\text{Pseudo}}$ (9)
Figure 5: EasyOcc model architecture: Image features are extracted and then processed in voxel space by 3D convolutions prior to pseudo-loss computation against our 3D pseudo-labels, Easy3D-Labels.
IV-C EASYOCC
Integrating $\mathcal{L}_{\text{Pseudo}}$ in EasyOcc is straightforward, as the framework relies exclusively on this signal for learning. The continuous and dense scene representation enables effective learning in conjunction with our 3D pseudo-labels.
The model is a simplified variant of GaussianOcc, with its architectural flow illustrated in Figure 5. Multi-view camera images are processed through a ResNet-101 image encoder to extract high-level features, which have shown robust performance across BEV models [11] and semantic occupancy prediction frameworks [9]. Parameter-free bilinear sampling projects these features into 3D space, which are then passed through a 3D CNN to enhance spatial reasoning and produce semantic logits. EasyOcc eliminates the need for depth estimation, multi-frame consistency,and novel view synthesis, thus reducing training complexity and duration.
V RESULTS
In this section, we present the main results. Subsections V-A and V-B describe the dataset, metrics, and configurations, followed by evaluations of Easy3D-Labels for mIoU and RayIoU in Subsections V-C and V-D. We then provide an ablation study of EasyOcc in Subsection V-E and qualitative analysis in Subsection V-F.
V-A DATASET AND EVALUATION METRICS
We evaluate all models on mIoU and RayIoU on the Occ3D-nuScenes dataset [45], and on RayIoU only on the OpenOccv2 dataset [46], as it does not contain a camera mask, which permits fair mIoU evaluation for self-supervised models. However, OpenOccv2 provides denser labels, which leads to a fairer evaluation on the RayIoU metric. Both datasets consist of 600 training scenes and 150 validation scenes from the nuScenes dataset. The voxel space is bounded by $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$, with a voxel size of $0.4\text{m}^{3}$.
For evaluation metrics, we utilize Intersection over Union (IoU), mean Intersection over Union (mIoU), and RayIoU. IoU, defined in Equation (10), reflects the model’s ability to capture overall spatial structure through occupancy. The mIoU metric, shown in Equation (11), computes the average IoU across all semantic classes, excluding the empty class.
$\text{IoU}=\frac{TP}{TP+FP+FN}$ (10)
$\text{mIoU}=\frac{1}{C}\sum_{\begin{subarray}{c}c=1\end{subarray}}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (11)
TP: True Positive, FP: False Positive, FN: False Negative
RayIoU is defined similarly to mIoU, but instead of being a voxel-wise metric, it is a ray metric. Introduced in SparseOcc [44], RayIoU aims to resolve the harsh penalisation of incorrect depth estimations and the overcompensation of overprediction of voxels for inflating the mIoU score. A ray is cast from the LiDAR sensor position in both the ground-truth and predicted voxel gird and it is labelled correct if both ray depths are within a threshold and are the same class. The equation is:
$\text{RayIoU}=\frac{1}{C}\sum_{\begin{subarray}{c}c=1\end{subarray}}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (12)
V-B MODEL CONFIGURATIONS
In Table I, we compare all model configurations. Rendering time and depth estimation add overhead, with Gaussian Splatting being the most efficient due to its rasterization-based rendering [21]. EasyOcc avoids these rendering methods during training, resulting in reduced training time. While incorporating pseudo-loss increases epoch training time, the impact varies by model, with OccNeRF experiencing the largest increase of (+31m) due to its overall slower training.
TABLE II: EasyOcc inference time breakdown.
Process Execution Time
Image Encoding 27ms
Bilinear Sampling 6ms
3D CNN 7ms
Grid Sampling 145ms
Total 185ms
TABLE III: State-of-the-art comparison on the Occ3D-nuScenes [45] dataset: FPS denotes frames per second, indicating processing time per sample. IoU and mIoU refer to Intersection over Union and mean Intersection over Union. Grey rows indicate SOTA methods for reference. The best result for each model (compared to its variant with our labels) is highlighted in bold. * OccNeRF uses 2D semantic loss, while GaussianOcc does not.
Method FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation DistillNeRF [48] 2.8 29.1 10.1 1.4 2.1 10.2 10.1 2.6 2.0 5.5 4.6 1.4 7.9 43.0 16.9 15.0 14.1 15.1 GaussTR [19] 0.3 44.5 13.8 6.5 8.5 21.8 24.3 6.3 15.5 7.9 1.9 6.1 17.2 37.0 17.2 7.2 21.2 10.0 TT-Occ [57] 0.7 - 16.7 21.5 10.5 10.7 14.7 11.9 12.3 9.7 12.2 4.4 7.9 48.3 23.7 28.3 14.1 20.2 GaussianFlowOcc [5] 10.2 46.9 17.1 6.8 9.7 19.0 17.2 4.2 11.8 9.3 10.3 1.8 12.3 61.0 31.2 34.8 14.7 12.4 AutoOcc [63] - 83.0 20.9 12.7 10.5 7.8 20.4 5.8 17.6 18.5 24.3 4.2 12.9 55.5 24.2 27.1 35.6 36.6 SelfOcc [17] 7.4 44.1 10.3 0.2 0.5 6.7 10.4 0.0 0.1 2.1 0.0 0.0 7.7 56.1 26.9 25.7 13.4 4.6 + Ours 7.4 34.5 14.4 1.7 5.4 14.5 22.2 2.6 6.4 15.4 8.9 1.0 12.8 55.0 26.8 21.6 11.3 9.3 OccNeRF [55] 5.4 46.4 11.0 0.7 1.8 6.6 6.6 3.7 0.3 2.9 3.2 2.9 6.6 52.8 24.0 25.0 18.6 9.7 + Ours 5.4 38.5 16.0 1.9 8.2 16.7 22.1 1.0 7.7 14.7 12.8 1.0 13.8 55.9 28.0 22.7 15.8 17.2 GaussianOcc [9] 5.4 42.9 11.3 1.8 5.8 14.6 13.6 1.3 2.8 8.0 9.8 0.6 9.6 44.6 20.1 17.6 8.6 10.3 + Ours 5.4 38.8 15.7 1.7 5.9 16.2 22.3 2.3 8.4 15.7 10.2 1.0 13.4 55.3 27.4 23.4 16.0 17.0 EasyOcc (Ours) 5.4 38.9 15.7 1.9 6.7 15.1 21.7 2.7 8.1 15.3 11.1 1.4 12.8 55.8 27.9 22.1 16.1 17.3
In Table II, we report the inference time (in milliseconds) for each component of the EasyOcc model. Grid Sampling refers to downsampling the contracted coordinate voxel grid to align with the dimensions of the Occ3D ground-truth labels, a technique introduced in OccNeRF [55]. Preliminary experiments showed that retaining the contracted coordinate system, rather than modeling the scene in the Occ3D output space, improved performance.
Self-supervised models that utilize LiDAR during training or inference are excluded from our comparison [23, 57, 42, 61] as our method is a camera-only pipeline, consistent with the models used for comparison in this study. Training and inference are performed on four NVIDIA A100-SXM4-40GB GPUs.
V-C MAIN RESULTS: MIOU
In this section, we evaluate the performance of the three selected baseline models and EasyOcc in Table III. We compare the models across four key evaluation categories: inference time (FPS), Intersection over Union (IoU), mean IoU (mIoU), and class-wise IoU for each semantic category. We provide further models for reference in the table.
V-C1 INFERENCE TIME
GaussianFlowOcc achieves the highest inference speed at 10.2 FPS, attributed to its use of induced attention, which significantly reduces computational overhead. SelfOcc ranks second with 7.4 FPS, benefiting from the lack of a contracted coordinate system. The inclusion of pseudo-loss has no effect on inference time, as it influences only the training phase. EasyOcc matches the inference speed of both OccNeRF and GaussianOcc, all of which operate at 5.4 FPS. Both GaussTR and TT-Occ exhibit poor FPS due to the deployment of foundation models at inference time.
V-C2 INTERSECTION OVER UNION
AutoOcc secures the highest performance with an IoU of 83.01, significantly surpassing all other models. GaussianFlowOcc ranks second, achieving an IoU of 46.9. Notably, the addition of pseudo-loss led to a marked decline in IoU; for example, it reduces SelfOcc’s score from 44.1 to 34.5, a 22% drop. This performance drop is attributed to object duplication, as the model lacks the ability to reason about occluded regions. Consequently, it tends to predict occupancy beyond visible surfaces, resulting in an overly dense scene representation, which is further evidenced in the evaluation of RayIoU in Subsection V-D. This is visualized in the qualitative analysis presented in Subsection V-F.
V-C3 MEAN INTERSECTION OVER UNION
A notable performance gap exists between the original voxel-based models (SelfOcc, OccNeRF, and GaussianOcc) and the Gaussian-based models: GaussTR, TT-Occ, GaussianFlowOcc, and AutoOcc. AutoOcc once again leads with an mIoU of 20.9, largely due to the integration of a VLM [8] and a VFM [60] during inference, which provides high-quality semantic estimations.
With the sole usage of Easy3D-Labels, EasyOcc reaches an mIoU of 15.7, surpassing even the Gaussian-based methods GaussTR. The employment of pseudo-loss allows OccNeRF to achieve a 15% improvement over GaussTR and a 45% gain compared to the original OccNeRF model. Similar performance boosts are observed for both SelfOcc and GaussianOcc. Notably, SelfOcc gains the ability to predict previously unsupported classes, such as construction vehicles, thanks to semantic supervision from Grounded-SAM. Furthermore, despite the lack of usage of rendering for loss, EasyOcc achieves results on par with the other three models, which use Easy3D-Labels, displaying the strength of our labels.
TABLE IV: RayIoU evaluated on the Occ3D-nuScenes [45] dataset: * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not. The best-performer is highlighted in bold, for each model compared to the same model trained with our labels integrated.
Method RayIoU RayIoU@1 RayIoU@2 RayIoU@4 $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [17] 9.6 6.7 9.6 12.5 0.7 1.0 15.9 16.3 0.0 0.9 5.0 0.0 0.0 20.8 48.7 13.4 17.2 15.9 8.0 + Ours 14.5 10.9 14.6 17.9 3.1 4.9 24.1 29.7 3.4 5.6 23.3 13.7 1.0 27.2 49.1 17.2 13.7 15.2 17.8 OccNeRF [55] 10.4 6.9 10.3 14.1 2.4 2.2 28.4 21.5 4.4 1.0 5.6 7.5 0.6 16.8 38.8 11.8 10.3 13.0 10.0 + Ours 15.5 11.7 15.6 19.2 3.2 6.5 27.2 30.0 3.2 6.5 23.5 16.2 1.9 29.5 50.2 18.0 14.6 15.1 20.2 GaussianOcc [9] 11.9 8.7 11.9 15.0 2.6 7.0 24.2 18.3 2.0 2.6 12.1 13.3 0.7 23.6 41.7 15.9 14.0 11.5 12.8 + Ours 14.3 10.7 14.4 17.8 2.9 3.8 23.4 29.4 3.6 6.2 21.8 12.3 0.8 25.7 48.9 17.4 13.5 15.0 20.0 EasyOcc (Ours) 14.6 10.9 14.7 18.2 3.1 4.9 22.7 28.7 4.8 6.6 21.4 13.2 2.4 27.4 49.8 17.2 13.3 14.8 19.9
TABLE V: RayIoU evaluated on the OpenOccv2 [46] dataset: * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not. The best-performer is highlighted in bold, for each model compared to the same model trained with our labels integrated.
Method RayIoU RayIoU@1 RayIoU@2 RayIoU@4 $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [17] 9.1 5.8 9.1 12.3 0.6 0.9 14.4 15.3 0.0 0.8 4.8 0.0 0.0 20.1 39.0 11.8 13.6 15.7 7.3 + Ours 15.1 11.2 15.3 18.7 2.9 4.3 23.2 28.5 3.5 5.1 22.7 12.4 1.3 26.6 45.4 16.3 18.0 16.2 19.2 OccNeRF [55] 11.4 7.9 11.3 15.0 2.1 1.8 29.4 21.6 4.3 0.9 4.9 6.2 0.6 17.3 41.8 13.8 13.8 12.7 9.6 + Ours 15.9 12.0 16.1 19.6 2.9 5.3 25.8 28.3 3.4 5.9 22.5 14.3 2.0 28.9 46.5 17.4 18.5 15.4 21.0 GaussianOcc [9] 11.7 8.5 11.8 14.8 2.4 5.8 22.8 17.7 2.0 2.2 11.0 11.1 0.6 23.3 37.8 14.9 13.7 11.1 12.2 + Ours 14.9 11.1 15.1 18.5 2.6 3.1 22.4 28.0 4.0 5.8 20.8 10.7 1.1 24.9 47.0 17.3 18.1 15.2 20.7 EasyOcc (Ours) 15.1 11.3 15.4 18.7 2.9 4.3 21.6 27.3 5.0 6.0 20.6 11.4 3.1 26.5 46.7 17.1 18.2 15.0 20.7
V-C4 IOU PER SEMANTIC CLASS
For single-class mIoU, AutoOcc leads in 6 out of 15 classes, particularly smaller objects such as bicycle, motorcycle, and traffic cone, likely due to the use of foundation models during inference. GaussianFlowOcc ranks first in 3 classes, mainly large-scale categories like drivable surface, sidewalk, and terrain, likely benefiting from strong temporal modeling. GaussTR achieves top performance in 5 classes, especially dynamic objects such as buses and cars, despite not using a dedicated flow module or temporal modeling.
With the integration of Easy3D-Labels into SelfOcc, OccNeRF, and GaussianOcc, results increase significantly across nearly all categories. For GaussianOcc, IoU increases in 14 out of 15 classes, with the barrier class as the sole exception. Similar trends are seen in SelfOcc and OccNeRF. We observe that performance generally improves for dynamic classes, while some static classes, particularly large-area categories such as manmade, do not show the same gains, potentially due to overprediction aiding their segmentation in models without our labels. Notably, SelfOcc demonstrates a dramatic 624% increase in IoU for the pedestrian class, emphasizing the importance of 3D labels for accurately detecting vulnerable object categories. EasyOcc exhibits strong detection capabilities, achieving performance comparable to that of the modified models enhanced with pseudo-loss.
V-D MAIN RESULTS: RAYIOU
In this section, we analyse models on the RayIoU metric across varying distance thresholds, along with class-wise RayIoU in Tables IV and V. As discussed previously, this metric penalises overprediction, which can otherwise inflate mIoU without accurately reflecting overall scene understanding. Additionally, evaluation across two datasets enables a more robust and consistent analysis.
Across both tables, we observe that training with Easy3D-Labels results in a substantial improvement in RayIoU for all distance thresholds. For example, OccNeRF shows a 49% and 39% increase on the Occ3D and OpenOccv2 datasets, respectively. Similar to the mIoU results, OccNeRF achieves the highest overall performance, reaching RayIoU scores of 15.5 and 15.9. Evaluating across different distance thresholds also highlights improved depth accuracy, with RayIoU@1 for SelfOcc increasing by over 93% on the OpenOccv2 dataset. Furthermore, these results emphasise the overprediction issue present in the original models, where IoU scores were artificially inflated, and for which RayIoU penalizes.
For class-wise RayIoU, Easy3D-Labels again enable more accurate detection, consistent with trends observed in the mIoU results, particularly for dynamic objects. Vulnerable classes, such as pedestrians, show significant improvements; for example, OccNeRF on Occ3D achieves an increase of over 300%, largely due to the accuracy provided by Metric3Dv2 [14]. Additionally, SelfOcc on OpenOccv2 shows improvements across all classes, with notable gains for car, motorcycle, and pedestrian. Our model, EasyOcc, demonstrates strong and competitive performance compared to SelfOcc trained with Easy3D-Labels, despite not relying on rendering-based losses.
V-E ABLATION STUDY: EASYOCC
In this section, we present a series of ablation studies on the EasyOcc model to assess the impact of key design choices. We ablate the image encoder, temporal samples, and pseudo-loss, among others. Models are trained for 12 epochs for computational efficiency.
V-E1 IMAGE ENCODER AND IMAGE SIZE
In Table VI, we investigate the effect of varying the image encoder and input image resolution on the mIoU metric, total model parameters, and inference speed (FPS).
TABLE VI: Image encoder and input image size ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
Encoder Image Size mIoU Model Parameters FPS
RN-152 15.3 56.5M 5.1
RN-101 14.9 40.9M 5.4
[…]
RN-34 14.9 16.2M 5.8
RN-18 14.4 10.8M 6.0
RN-152 13.8 56.5M 5.4
RN-101 13.6 40.9M 5.6
[…]
RN-34 13.2 16.2M 6.0
RN-18 12.8 10.8M 6.2
[…]
V-E2 AGGREGATION OF TEMPORAL SAMPLES
As shown in the previous Figure 2, the use of temporal samples significantly improves the similarity of our 3D pseudo-labels to the Occ3D ground-truth annotations. In Figure 6, we extend this analysis by training the EasyOcc model with varying numbers of temporal samples to evaluate whether a similar trend holds in model performance post-training.
Figure 6: Temporal sample ablation.
The figure exhibits a logarithmic curve, seen previously when comparing Occ3D ground-truth labels to the 3D pseudo-labels, with the results clearly showing that increasing the number of temporal samples improves the mIoU metric. This improvement is attributed to scene densification, especially in regions farther from the camera’s field of view. Notably, we achieve an mIoU of 14.9 with both 10 and 13 temporal samples, further indicating a saturation point beyond which additional samples yield diminishing returns in the mIoU metric.
V-E3 VARIATION OF $\lambda$
In Table VII, we experiment with the $\lambda$ constant in Equation (5) to balance the contributions of the cross-entropy and geometry loss components. The results show that a $\lambda$ value of 0.1 yields the best mIoU performance. A value of 1 performs slightly worse, trailing by 0.2 points, while a value of 0.01 results in a significant drop of 0.7 points. These findings indicate that geometry loss plays a meaningful role, but balancing the contributions of both cross-entropy and geometry losses is crucial. With $\lambda=0.1$, the magnitudes of the cross-entropy and geometry losses are approximately equal during testing, underscoring the importance of careful weighting for optimal performance.
TABLE VII: Pseudo-loss $\lambda$ ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
Lambda $\lambda$ mIoU
0.01 14.0
0.1 14.9
1 14.7
V-E4 CHOICE OF LOSSES
In Table VIII, we ablate the pseudo-loss in Equation (5) by removing individual loss components in Equation (6), to assess the contribution of each component to the model’s learning.
TABLE VIII: Pseudo-loss ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
Cross En. Geometry Scale Semantic Scale Lovász mIoU
14.9
15.1
13.8
14.2
12.5
9.2
9.9
Overall, the results show that removing components of the loss function reduces the model’s learning capability. For example, excluding the cross-entropy loss decreases mIoU by 1.1 points. Similar drops occur when individual components of the geometry loss are removed. Interestingly, excluding the Lovász-Softmax loss results in a slight mIoU improvement of 0.2, which is unexpected since it is designed to optimize the IoU metric. We hypothesize that this anomaly arises from including the empty class index in the Lovász loss computation. To validate this, we retrain the model excluding this index, as detailed in the following ablation.
V-E5 LOVÀSZ SOFTMAX LOSS
Following the previous experiment, we retrain the model with the Lovász-Softmax loss that excludes the empty index from the loss computation, seen in Table IX.
TABLE IX: Lovász-Softmax empty index ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
Ignore Empty mIoU
14.9
15.4
Excluding the empty class from the loss computation results in a 0.5 mIoU improvement. Only the previous experiments in the ablation section were conducted with the inclusion of the empty label in the Lovász loss. However, the integrity of the comparative results is maintained, as all models were subject to the same conditions and limitations.
V-F QUALITATIVE RESULTS
In this section, we present a qualitative analysis in two parts. First, we compare predicted semantic voxels against ground-truth labels. Then, we compare voxel estimations across the six camera views for GaussianOcc and its variant trained with our pseudo-loss to assess the benefits of Easy3D-Labels.
V-F1 VOXEL ANALYSIS
Figure 7: Voxel qualitative analysis of models evaluated in Table III on sample token e67f3e81225f426f8e1743af45487762. +(Ours) indicates the same model trained with Easy3D-Labels.
We begin by analysing Figure 7, which shows semantic voxel visualisations for each model, comparing original models with their pseudo-loss variants.
SelfOcc performs well in road segmentation but misses vehicles in the back camera view; pseudo-loss corrects this but introduces increased misclassification as barrier, particularly in the back-left view, likely due to confusing structures such as fences or signs.
OccNeRF predicts more correct objects than SelfOcc, including vehicles in the back view, but suffers from object duplication, leading to penalties in metrics. Pseudo-loss reduces duplication and improves vegetation estimation. Similar improvements are observed in GaussianOcc, where pseudo-loss corrects misclassifications such as road being predicted as a wall.
Comparing EasyOcc and GaussTR, GaussTR fails to capture vegetation and building overhangs, while EasyOcc predicts these correctly. However, GaussTR better reconstructs dynamic objects, producing more complete shapes, whereas EasyOcc predicts only visible regions. This aligns with quantitative results, where GaussTR performs better on dynamic classes. EasyOcc, using a voxel representation, produces smoother outputs, while GaussTR’s Gaussian representation appears more fragmented.
Easy3D-Labels provide several benefits: they enable correct estimations of regions beneath the ego vehicle through temporal aggregation, improve detection of structures such as building overhangs absent in ground truth, and reduce object duplication and scene densification, leading to improved mIoU and RayIoU.
V-F2 IMAGE VIEW ANALYSIS
Figure 8: Image view qualitative analysis of models evaluated in Table III on sample token e67f3e81225f426f8e1743af45487762. +(Ours) indicates the same model trained with Easy3D-Labels.
Model estimations and our 3D pseudo-labels are generated solely from camera views. We therefore examine semantic voxel estimations of GaussianOcc and its variant trained with Easy3D-Labels across the six camera views in Figure 8.
In the front left view, a ground-truth mislabel causes GaussianOcc to be penalised despite correctly predicting a pole, while our model predicts vegetation. In the front right view, both models detect a pedestrian but with inaccurate positioning due to occlusion; neither correctly labels smaller objects such as the fire hydrant or trash can, though both detect the motorcycle.
In the back right view, our model correctly captures a building overhang, which GaussianOcc misses. In the back view, GaussianOcc fails to predict the road in the lower region and incorrectly introduces a wall behind vehicles, while our model produces a continuous road and avoids this error. In the back left view, GaussianOcc generates several unsupported estimations (e.g., barrier, pedestrian, construction vehicle), whereas our model produces cleaner and more consistent outputs.
VI CONCLUSION
This paper presents the use of 3D pseudo-labels, Easy3D-Labels, for self-supervised semantic occupancy estimation models for automated vehicle perception. These labels enable loss computation directly in 3D space, rather than relying on conventional 2D camera-space supervision. They can be easily integrated into existing architectures, leading to improved model performance, more complete scene representation, and better detection of vulnerable road users. Additionally, using only these labels for supervision in our model, EasyOcc, proves effective across performance metrics. The strong performance of these 3D pseudo-labels highlights their potential to enhance self-supervised models. However, several directions for future work remain to further refine and evaluate the proposed approach:
  1. 1.
    Incorporating LiDAR data into the 3D pseudo-label generation pipeline to facilitate comparison with LiDAR-supervised models.
  2. 2.
    Conducting a more comprehensive investigation into the integration of 3D pseudo-labels within models that utilise a Gaussian scene representation.
  3. 3.
    Evaluating the robustness of 3D pseudo-labels under challenging driving conditions, such as rain, fog, and low-light environments.
Self-supervised occupancy estimation models have historically lagged behind supervised methods, but recent advancements, including this work, indicate they are closing the performance gap. While research is shifting toward Gaussian representations, the adoption of 3D pseudo-labels for voxel-based models remains uncertain. Nonetheless, this study shows that leveraging temporal information and carefully selecting the loss domain are key to achieving strong performance in semantic occupancy estimation.
Acknowledgement
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.